At a checkout counter customers arrive at an average of 2 per minute. Find the
probability that
(a) at most 4 will arrive at any given minute.
(b) at least 3 will arrive during an interval of 2 minutes
(c) 5 will arrive in an interval of 3 minutes.
1
Expert's answer
2020-04-27T19:27:31-0400

Assume that customers arrival has Poisson distribution with λ=2

Let Pk=the number of arrivals during the given interval.

a) λ=2

P(k≤4)=P0+P1+P2+P3+P4= e-λ0/0!+λ1/1!+λ2/2!+λ3/3!+λ4/4!)=0.1429.

b) λ=2*2=4

P(k≥3)=1−P0​−P1​−P2​=1−(40/0!)​⋅e−4−(41/1!)​⋅e−4-(42/2!)​⋅e−4=1-e-4(1+4+8)=0.7619.

с) Poisson distribution with λ=2*3=6.

P(X=5)=e55/5!)​=0.1606.